Find the effective interest rate for an investment that earns interest of 10% per annum if interest compounds I) yearly II) monthly iii) half yearly iv) bi-monthly
1
Expert's answer
2020-12-13T17:17:02-0500

1)

"r=((1+\frac{10\%}{1\cdot100\%})^1-1)\cdot100\%=10.00\%."

2)

"r=((1+\frac{10\%}{12\cdot100\%})^{12}-1)\cdot100\%=10.47\%."

3)

"r=((1+\frac{10\%}{2\cdot100\%})^2-1)\cdot100\%=10.25\%."

4)

"r=((1+\frac{10\%}{24\cdot100\%})^{24}-1)\cdot100\%=10.49\%."